# Is uniform expressivity too restrictive? Towards efficient expressivity of GNNs

### v1

@@ Appendix C Proofs for Section 5
Let us overview the proof of Theorem 5.1. First, we will show that the constructive proof of (Barceló et al., 2020, Proposition 4.2) for $\operatorname{CReLU}$ holds for the linear threshold activation function $\sigma_{*}$ defined in equation 1. After this, we will exploit that for a fast convergence of step-like function (see Definition 5.1) $\sigma$, $\sigma^{m}$ converges fast to $\sigma_{*}$ as the following lemma shows.
@@
Let $\sigma$ be a step-like activation function. Then for all $x\notin(\varepsilon,1-\varepsilon)$ and all $m\geq N$,
$\left|\sigma^{m}(x)-\sigma_{*}(x)\right|\leq\varepsilon\left(\frac{1+\eta}{2}\right)^{m-N}.$ (3)
Moreover, if further $\eta=0$, then for all $x\notin(\varepsilon,1-\varepsilon)$ and all $m\geq N$,
$\left|\sigma^{m}(x)-\sigma_{*}(x)\right|\leq\frac{\varepsilon}{2^{2^{m-N}-1}}.$ (4)
@@
Let $Q$ be a query of GC2 of depth $d$, and let $(Q_{1},\cdots,Q_{d})$ be an enumeration of its subformulas. First, we extend the constructive proof of (Barceló et al., 2020, Proposition 4.2) to GNNs with Linear Threshold activations $\sigma_{*}$. In this extension, we have to make sure we can control the error made efficiently [^4] The original proof to exactly compute the query can actually be extended for any function $g:\mathbb{R}\rightarrow\mathbb{R}$ verifying: there exists real $a<b$ such that i) for every real $x\leq a,g(x)=0$, ii) for every real $x\geq b,g(x)=1$ and iii) $g$ is non-decreasing on $[a,b]$. .
We consider the combination functions for $t\in[d]$, $\mathsf{comb}_{t}:(x,y)\mapsto\sigma_{*}(A_{t}x+B_{t}y+C_{t})$, with the same choices of $d\times d$ matrices $A_{t}$, $B_{t}$ and $C_{t}$ initially set to 00 and filled as described below. Each row of $A_{t}$, $B_{t}$ and $C_{t}$ corresponds to a subformula $Q_{j}$ with $j\leq t$ as follows:
1) (Queries of depth $1$) The $\ell^{\prime}\leq d$ color functions used by $Q_{j}$ are encoded in the first $\ell^{\prime}$ rows of $A_{1}$. To do so, the first $\ell^{\prime}$ rows are the transpose of the first $i$-th canonical vectors of $\mathbb{R}^{d}$ for $i\in[\ell^{\prime}]$.
2) For $j>1$, if $Q_{j}(x)=\exists^{\geq K}yE(x,y)\land Q_{i}(y)$ for some $i$ with $j>i$. Set $C_{j}=-(K-1)$, $A_{j,k}=0$ for every $k\in[j]$, $B_{j,i}=1$ and $B_{j,k}=0$ for every $k\in[j]$.
3) For $j>1$, if $Q_{j}=\lnot Q_{i}$ for some $i$ with $j>i$. Set $C_{j}=+1$, $A_{j,i}=-1$ and $A_{j,k}=0$ for $k\in[j]-\{i\}$, and $B_{j,k}=0$ for every $k\in[j]$.
4) For $j>2$, if $Q_{j}=Q_{i_{1}}\land Q_{i_{2}}$ for some $i_{1},i_{2}$ in $[j-1]$. Set $C_{j}=-1$, $A_{j,i_{1}}=A_{j,i_{2}}=1$, $A_{j,k}=0$ for $k\in[j]-\{i_{1},i_{2}\}$, and $B_{j,k}=0$ for every $k\in[j]$.
For any integer $t<d$ and for every graph $G$ and each vertex $v\in V(G)$ of $G$, consider
$q^{t+1}(G,v):=\mathsf{comb}_{t}\left(q^{t}(G,v),\sum_{w\in\mathcal{N}_{G}(v)}q^{t}(G,w)\right).$ (5)
@@
Observe that for each $t\in[d]$, $q^{t}(G,v)\in\{0,1\}^{d}$. With the exact same inductive argument as in (Barceló et al., 2020, Proposition 4.2), we have that
$q^{t}(G,v)=(Q_{1}(G,v),\ldots,Q_{d}(G,v))^{T}.$
@@
We will now consider the GNN where we have substituted $\sigma_{*}$ by its approximation $\sigma^{m}$. The new GNN will have combination functions of the form $\mathsf{comb}_{t}^{m}:(x,y)\mapsto\sigma^{m}(A_{t}x+B_{t}y+C_{t})$, where $t\in[d]$, and $\xi^{m}_{t+1}(G,v)=\mathsf{comb}_{t}^{m}\left(\xi^{m}_{t}(G,v),\sum_{w\in\mathcal{N}_{G}(v)}\xi^{m}_{t}(G,w)\right)$, where $t\in[d-1]$. Observe that $\mathsf{comb}_{t}^{m}$ is a neural network of width $1$ and depth $m$ with activation function $\sigma$.
@@
Let $\Delta$ be a positive integer and fix a graph $G$ with maximal degree bounded by $\Delta$. For $t\in[d]$, we define
$\epsilon_{t}:=\max_{v\in V(G)}\|\xi^{t}(G,v)-q^{t}(G,v)\|_{\infty}.$
@@
Claim. For any $t\in[d-1]$,
if $\text{if }2(\Delta+2)\epsilon_{t}<1\text{, then }\epsilon_{t+1}\leq\begin{cases}\left(\frac{1+\eta}{2}\right)^{m-N}\varepsilon&\text{if }\eta>0\\ \frac{\varepsilon}{2^{2^{m-N}-1}}&\text{if }\eta=0\end{cases}$
@@
$\eta>0\;\text{and}\;m\geq N+\frac{2+\log(\Delta+2)}{1-\log(1-\eta)}$
@@
$2(\Delta+2)\epsilon_{t}\implies 2(\Delta+2)\epsilon_{t+1}$
@@
$\epsilon_{d}\leq\begin{cases}\left(\frac{1+\eta}{2}\right)^{m-N}\varepsilon&\text{if }\eta>0\\ \frac{\varepsilon}{2^{2^{m-N}-1}}&\text{if }\eta=0\end{cases}\leq 1/3<1/2,$
@@
Now, we prove the claim. For $t\in[d-1]$ we have by the triangular inequality
$\epsilon_{t+1}$ $=\max_{v\in V(G)}\|\xi^{t+1}(G,v)-q^{t+1}(G,v)\|_{\infty}$
$=\max_{v\in V(G)}\left\|\mathsf{comb}_{t}^{m}\left(\xi^{t}(G,v),\sum_{w\in\mathcal{N}_{G}(v)}\xi^{t}(G,w)\right)-\mathsf{comb}_{t}\left(q^{t}(G,v),\sum_{w\in\mathcal{N}_{G}(v)}q^{t}(G,w)\right)\right\|_{\infty}$
$\leq\max_{v\in V(G)}\left\|\mathsf{comb}_{t}^{m}\left(\xi^{t}(G,v),\sum_{w\in\mathcal{N}_{G}(v)}\xi^{t}(G,w)\right)-\mathsf{comb}_{t}\left(\xi^{t}(G,v),\sum_{w\in\mathcal{N}_{G}(v)}\xi^{t}(G,w)\right)\right\|_{\infty}$
$\quad+\max_{v\in V(G)}\left\|\mathsf{comb}_{t}\left(\xi^{t}(G,v),\sum_{w\in\mathcal{N}_{G}(v)}\xi^{t}(G,w)\right)-\mathsf{comb}_{t}\left(q^{t}(G,v),\sum_{w\in\mathcal{N}_{G}(v)}q^{t}(G,w)\right)\right\|_{\infty}.$
@@
$\sup\left\{\left\|\mathsf{comb}_{t}^{m}\left(x,y\right)-\mathsf{comb}_{t}\left(x,y\right)\right\|_{\infty}:\;\text{for all}\;i,\,x_{i}\notin(\epsilon_{t},1-\epsilon_{t}),\,y_{i}\notin(\Delta\epsilon_{t},1-\Delta\epsilon_{t})\right\},$
since each component of $\xi^{t}(G,v)$ is at distance at most $\epsilon_{t}$ from 00 or $1$ by definition of $\epsilon_{t}$. Now, $\mathsf{comb}_{t}^{m}\left(x,y\right)=\sigma^{m}(A_{t}x+B_{t}x+C_{t})$ and $\mathsf{comb}_{t}\left(x,y\right)=\sigma_{*}(A_{t}x+B_{t}y+C_{t})$ with $A_{t},B_{t}\in\{-1,0,1\}^{d\times d}$ and $C_{t}\in\mathbb{Z}^{d}$. Moreover, the matrix $A_{t}$ has at most two non-zero entries per row, and $B_{t}$ at most one non-zero entry. Since the $1$-norm of each row of $A_{t}$ is at most $2$, and the one of $B_{t}$ is at most $1$, we can improve the upper-bound with
$\sup\left\{|\sigma^{m}\left(z\right)-\sigma_{*}\left(z\right)|:z\notin((\Delta+2)\epsilon_{t},1-(\Delta+2)\epsilon_{t})\right\},$
@@
$x^{\prime}:=q^{t}(G,v)~{}\text{ and }y^{\prime}:=\sum_{w\in\mathcal{N}_{G}(v)}q^{t}(G,w)$
@@
$\left\|(A_{t}x+B_{t}y+C_{t})-(A_{t}x^{\prime}+B_{t}y^{\prime}+C_{t})\right\|_{\infty}=\left\|A_{t}(x-x^{\prime})+B_{t}(y-y^{\prime})\right\|_{\infty}\leq(2+\Delta)\epsilon_{t}.$
In this way, $A_{t}x+B_{t}y+C_{t}$ is a real vector with $\infty$-norm $<1/2$ to the integer vector $A_{t}x^{\prime}+B_{t}y^{\prime}+C_{t}$ and so we conclude that
$\mathsf{comb}_{t}(x,y)=\mathsf{comb}_{t}(x^{\prime},y^{\prime}),$
because $\sigma_{*}$ takes the same value on an integer and on a real number at distance less than $\frac{1}{2}$ from it. Hence the value of the difference under the second maximum is zero, proving the claim. ∎
@@
$|\sigma^{m}(\sigma^{N}(x))-\sigma_{*}(x)|$
@@
This is a combination of Theorem 5.1 with Proposition 5.2. ∎
@@
This is a combination of Theorem 5.1 with Proposition 5.3. ∎
@@
This is a combination of Theorem 5.1 with Proposition 5.4. ∎

### v2

@@ Appendix C Proofs for Section 5
Let us overview the proof of Theorem 5.1. First, we will show that the constructive proof of (Barceló et al., 2020, Proposition 4.2) for $\operatorname{CReLU}$ holds for the linear threshold activation function $\sigma_{*}$ defined in equation 2. After this, we will exploit that for a fast convergence of step-like function (see Definition 5.1) $\sigma$, $\sigma^{\ell}$ converges fast to $\sigma_{*}$ as the following lemma shows.
@@
Let $\sigma$ be a step-like activation function. Then for all $x\notin(\varepsilon,1-\varepsilon)$ and all $\ell\geq N$,
$\left|\sigma^{\ell}(x)-\sigma_{*}(x)\right|\leq\varepsilon\left(\frac{1+\eta}{2}\right)^{\ell-N}.$ (4)
Moreover, if further $\eta=0$, then for all $x\notin(\varepsilon,1-\varepsilon)$ and all $\ell\geq N$,
$\left|\sigma^{\ell}(x)-\sigma_{*}(x)\right|\leq\frac{\varepsilon}{2^{2^{\ell-N}-1}}.$ (5)
@@
Let $Q$ be a query of GC2 of depth $d$, and let $(Q_{1},\ldots,Q_{d})$ be an enumeration of its subformulas, being each $Q_{i}$ an atomic formula or a formula build from the previous ones. Without loss of generality, assume that $Q$ uses all $\ell\leq d$ colors and that the first $\ell$ subformulas $Q_{1},\ldots,Q_{\ell}$ are the atomic formulas
$\mathsf{Col}_{1}(x),\ldots,\mathsf{Col}_{\ell}(x)$
corresponding to these $\ell$ colors. If this is not the case, then $Q$ involves $\ell^{\prime}\leq\ell$ colors, say $\lambda_{1},\ldots,\lambda_{\ell^{\prime}}$, and so in the sequence $(Q_{1},\ldots,Q_{d})$, we will only have the atomic formulas $\mathsf{Col}_{\lambda_{1}}(x),\ldots,\mathsf{Col}_{\lambda_{\ell^{\prime}}}(x)$ appearing. Thus to make the construction below work, we only need to reorder the subformulas and to perform the constructions below changing the color $\lambda_{i}$ to the color $i$ in the formulas and change the $A$ and $B$ in the first combination function $\mathsf{comb}_{0}$ from $A$ and $B$ to $AP$ and $BP$ where $P$ is the coordinate projection mapping $e_{\lambda_{i}}$ to $e_{i}$.
@@
First, we extend the constructive proof of (Barceló et al., 2020, Proposition 4.2) to GNNs with linear threshold activations $\sigma_{*}$. In this extension, we have to make sure we can control the error made efficiently [^3] The original proof to exactly compute the query can actually be extended for any function $g:\mathbb{R}\rightarrow\mathbb{R}$ verifying: there exists real $a<b$ such that i) for every real $x\leq a,g(x)=0$, ii) for every real $x\geq b,g(x)=1$ and iii) $g$ is non-decreasing on $[a,b]$.. The GNN that we will build has $d$ iterations, being each combination functions of the form
$\mathsf{comb}_{t}:(x,y)\mapsto\sigma_{*}(Ax+By+C)\qquad(t\in\{0,\ldots,d-1\})$
where $A$ and $B$ are $d\times d$ matrices and $C$ a $d$-dimensional vector $C$. The $j$th row of $A$, $B$ and $C$ will correspond to the corresponding subformula $Q_{j}$ and they will be filled as follows:
1) If $Q_{j}$ is the atomic formula corresponding to the color $\lambda_{j}$, then $A_{j,\lambda_{j}}=1$ for $k=\lambda_{j}$ and $A_{j,k}=0$ otherwise, $B_{j,k}=0$ for all $k\in[d]$, and $C_{j}=0$. Note that by our assumption, the first $\ell$ rows of $A$ are those of the identity matrix, and so the first $\ell$ components of the embeddings give the color of the vertex.
2) If $Q_{j}(x)=\exists^{\geq K}y\left(E(x,y)\land Q_{i}(y)\right)$ for some $i<j$, then $A_{j,k}=0$ for all $k\in[d]$, $B_{j,i}=1$ and $B_{j,k}=0$ for $k\neq i$, and $C_{j}=-(K-1)$.
3) If $Q_{j}(x)=\exists^{\geq K}y\left(E(x,y)\land\textsc{t}\right)$, then $A_{j,k}=0$ for all $k\in[d]$, $B_{j,k}=1$ for $k\in[\ell]$ and $B_{j,k}=0$ for $k\notin[\ell]$, and $C_{j}=-(K-1)$.
4) If $Q_{j}(x)=\lnot Q_{i}(x)$ with $i<j$, then $A_{j,i}=-1$ and $A_{j,k}=0$ for $k\neq i$, $B_{j,k}=0$ for all $k\in[d]$ and $C_{j}=1$.
5) If $Q_{j}(x)=\lnot\textsc{t}$, then $A_{j,i}=B_{j,i}=0$ for all $k\in[d]$ and $C_{j}=0$.
6) If $Q_{j}(x)=Q_{i_{1}}(x)\land Q_{i_{2}}(x)$ for some $i_{1},i_{2}<j$, then $A_{j,i_{1}}=A_{j,i_{2}}=1$ and $A_{j,k}=0$ for $k\neq i_{1},i_{2}$, $B_{j,k}=0$ for all $j\in[d]$ and $C_{j}=-1$.
7) If $Q_{j}(x)=Q_{i}(x)\land\textsc{t}$ or $Q_{j}(x)=\textsc{t}\land Q_{i}(x)$ for some $i<j$, then $A_{j,i}=1$ and $A_{j,k}=0$ for $k\neq i$, $B_{j,k}=0$ for all $j\in[d]$ and $C_{j}=0$.
8) If $Q_{j}(x)=\textsc{t}\land\textsc{t}$, then $A_{j,k}=B_{j,k}=0$ for all $j\in[d]$ and $C_{j}=1$.
For any integer $t\in\{0,\ldots,d-1\}$ and for every graph $G$ and each vertex $v\in V(G)$ of $G$, consider
$q^{t+1}(v,G):=\mathsf{comb}_{t}\left(q^{t}(v,G),\sum_{w\in\mathcal{N}_{G}(v)}q^{t}(w,G)\right).$ (6)
@@
Observe that for each $t\in[d]$, $q^{t}(v,G)\in\{0,1\}^{d}$. With the exact same inductive argument as in (Barceló et al., 2020, Proposition 4.2), we have that for all $t\in[d]$ and all $j\in[t]$,
$q^{t}_{j}(v,G)=Q_{j}(v,G),$
i.e., the first $t$ components of the embedding $q^{t}(v,G)$ are equal to the first $t$ subformulas $Q_{1},\ldots,Q_{t}$ of $Q$.
@@
Now, we will now consider the GNN where we have substituted $\sigma_{*}$ by its approximation $\sigma^{\ell}$. The new GNN will have combination functions of the form
$\mathsf{comb}_{t}^{\ell}:(x,y)\mapsto\sigma^{\ell}(Ax+By+C)\qquad(t\in\{0,\ldots,d-1)$
with $A$, $B$ and $C$ as above, and it will construct the vertex embedding $\xi^{\ell}_{t}(v,G)$ given by the update rule
$\xi^{\ell}_{t+1}(v,G)=\mathsf{comb}_{t}^{\ell}\left(\xi^{\ell}_{t}(v,G),\sum_{w\in\mathcal{N}_{G}(v)}\xi^{\ell}_{t}(w,G)\right).$
Observe that $\mathsf{comb}_{t}^{\ell}$ is a neural network with input size $2d$, output size $d$, $\ell+1$ layers and widths $2d,d,\ldots,d$, and so size $(\ell+2)d$.
@@
Let $\Delta$ be a positive integer and fix a graph $G$ with maximal degree bounded by $\Delta$. For $t\in\{0,\ldots,d\}$, we define
$\epsilon_{t}:=\max_{v\in V(G)}\|\xi^{t}(v,G)-q^{t}(v,G)\|_{\infty}.$
@@
Claim. For any $t\in\{0,\ldots,d-1\}$,
if $\text{if }2(\Delta+2)\epsilon_{t}<1\text{, then }\epsilon_{t+1}\leq\begin{cases}\left(\frac{1+\eta}{2}\right)^{\ell-N}\varepsilon&\text{if }\eta>0\\ \frac{\varepsilon}{2^{2^{\ell-N}-1}}&\text{if }\eta=0\end{cases}$
@@
$\eta>0\;\text{and}\;\ell\geq N+\frac{2+\log(\Delta+2)}{1-\log(1+\eta)}$
@@
$12(\Delta+2)\epsilon_{t}<1\implies 2(\Delta+2)\epsilon_{t+1}<$ 1
@@
$\epsilon_{d}\leq\begin{cases}\left(\frac{1+\eta}{2}\right)^{\ell-N}\varepsilon&\text{if }\eta>0\\ \frac{\varepsilon}{2^{2^{\ell-N}-1}}&\text{if }\eta=0\end{cases}\leq 1/3<1/2,$
@@
Now, we prove the claim. For $t\in\{0,\ldots,d-1\}$ we have, by the triangle inequality, that
$\epsilon_{t+1}$ $=\max_{v\in V(G)}\|\xi^{t+1}(v,G)-q^{t+1}(v,G)\|_{\infty}$
$=\max_{v\in V(G)}\left\|\mathsf{comb}_{t}^{\ell}\left(\xi^{t}(v,G),\sum_{w\in\mathcal{N}_{G}(v)}\xi^{t}(w,G)\right)-\mathsf{comb}_{t}\left(q^{t}(v,G),\sum_{w\in\mathcal{N}_{G}(v)}q^{t}(w,G)\right)\right\|_{\infty}$
$\leq\max_{v\in V(G)}\left\|\mathsf{comb}_{t}^{\ell}\left(\xi^{t}(v,G),\sum_{w\in\mathcal{N}_{G}(v)}\xi^{t}(w,G)\right)-\mathsf{comb}_{t}\left(\xi^{t}(v,G),\sum_{w\in\mathcal{N}_{G}(v)}\xi^{t}(w,G)\right)\right\|_{\infty}$
$\quad+\max_{v\in V(G)}\left\|\mathsf{comb}_{t}\left(\xi^{t}(v,G),\sum_{w\in\mathcal{N}_{G}(v)}\xi^{t}(w,G)\right)-\mathsf{comb}_{t}\left(q^{t}(v,G),\sum_{w\in\mathcal{N}_{G}(v)}q^{t}(w,G)\right)\right\|_{\infty}.$
@@
$\sup\left\{\left\|\mathsf{comb}_{t}^{\ell}\left(x,y\right)-\mathsf{comb}_{t}\left(x,y\right)\right\|_{\infty}:\;\text{for all}\;i,\,x_{i}\notin(\epsilon_{t},1-\epsilon_{t}),\,y_{i}\notin(\Delta\epsilon_{t},1-\Delta\epsilon_{t})\right\},$
since each component of $\xi^{t}(v,G)$ is at distance at most $\epsilon_{t}$ from 00 or $1$ by definition of $\epsilon_{t}$. Now, $\mathsf{comb}_{t}^{\ell}\left(x,y\right)=\sigma^{\ell}(Ax+Bx+C)$ and $\mathsf{comb}_{t}\left(x,y\right)=\sigma_{*}(Ax+By+C)$ with $A,B\in\{-1,0,1\}^{d\times d}$ and $C\in\mathbb{Z}^{d}$. Moreover, the matrix $A$ has at most two non-zero entries per row, and $B$ at most one non-zero entry. Since the $1$-norm of each row of $A$ is at most $2$, and the one of $B$ is at most $1$, we can improve the upper-bound with
$\sup\left\{|\sigma^{\ell}\left(z\right)-\sigma_{*}\left(z\right)|:z\notin((\Delta+2)\epsilon_{t},1-(\Delta+2)\epsilon_{t})\right\},$
@@
$x^{\prime}:=q^{t}(v,G)~{}\text{ and }y^{\prime}:=\sum_{w\in\mathcal{N}_{G}(v)}q^{t}(w,G)$
@@
$\left\|(Ax+By+C)-(Ax^{\prime}+By^{\prime}+C)\right\|_{\infty}=\left\|A(x-x^{\prime})+B(y-y^{\prime})\right\|_{\infty}\leq(2+\Delta)\epsilon_{t}.$
In this way, $Ax+By+C$ is a real vector with $\infty$-norm $<1/2$ to the integer vector $Ax^{\prime}+By^{\prime}+C$ and so we conclude that
$\mathsf{comb}(x,y)=\mathsf{comb}(x^{\prime},y^{\prime}),$
because $\sigma_{*}$ takes the same value on an integer and on a real number at distance less than $1/2$ from it. Hence the value of the difference under the second maximum is zero, proving the claim. ∎
@@
$|\sigma^{\ell}(\sigma^{N}(x))-\sigma_{*}(x)|$
@@
This is a combination of Theorem 5.1 with Proposition 5.2 using Proposition C.4 below. ∎
@@
This is a combination of Theorem 5.1 with Proposition 5.3 using Proposition C.4 below. ∎
@@
This is a combination of Theorem 5.1 with Proposition 5.4 below. ∎
Proposition C.4.
Let $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ and $\tilde{\sigma}:\mathbb{R}\rightarrow\mathbb{R}$ be functions. If $\tilde{\sigma}$ can be written in terms of a NN with activation function $\sigma$ with $\ell+2$ layers of widths $11,r_{1},\ldots,r_{\ell},$ 1, then every NN with activation function $\tilde{\sigma}$ with $k+1$ layers of widths $w_{0},\ldots,w_{k+1}$ can be transformed into a NN with activation function $\sigma$ with $2+\ell(k-1)$ layers of widths
$w_{0},w_{1}r_{1},\ldots,w_{1}r_{\ell},w_{2}r_{1},\ldots,w_{2}r_{\ell},\ldots,w_{k}r_{1},\ldots,w_{k}r_{\ell},w_{k+1}.$
In particular, if $\tilde{\sigma}$ can be written in terms of a NN with activation function $\sigma$ of size $r$, then every NN with activation function $\tilde{\sigma}$ of size $s$ can be transformed into a NN with activation function $\sigma$ of size at most $(r-2)s$.
Proof.
Let the NN with activation function $\sigma$ have affine transformations $T_{i}:\mathbb{R}^{w_{i}}\rightarrow\mathbb{R}^{w_{i+1}}$ ($i\in\{0,\ldots,k\}$) and the NN with activation function $\sigma$ that expresses $\tilde{\sigma}$ have affine transformations $S_{i}:\mathbb{R}^{r_{i}}\rightarrow\mathbb{R}^{r_{i+1}}$ ($i\in\{0,\ldots,\ell+1\}$), where $r_{0}=r_{\ell+2}=1$. Then, as an $\mathbb{R}\rightarrow\mathbb{R}$ map we have
$\tilde{\sigma}=S_{\ell+1}\circ\sigma\circ S_{\ell}\circ\cdots S_{1}\circ\sigma\circ S_{0},$
but, as a pointwise $\mathbb{R}^{w}\rightarrow\mathbb{R}^{w}$ map, we can write $\tilde{\sigma}$ as
$\tilde{\sigma}=S_{\ell+1}^{\otimes w}\circ\sigma\circ S_{\ell}^{\otimes w}\circ\cdots S_{1}\circ\sigma\circ S_{0}^{\otimes w}$
where $S_{i}^{\otimes w}$ is the map $\mathbb{R}^{r_{i}w}\rightarrow\mathbb{R}^{r_{i+1}w}$ given by
$\mathbb{R}^{r_{i}w}=(\mathbb{R}^{r_{i}})^{w}\ni\begin{pmatrix}x_{1}\\ \vdots\\ x_{w}\end{pmatrix}\mapsto\begin{pmatrix}S_{i}(x_{1})\\ \vdots\\ S_{i}(x_{w})\end{pmatrix}\in(\mathbb{R}^{r_{i+1}})^{w}=\mathbb{R}^{r_{i+1}w}.$
In this way, considering the NN with activation function $\sigma$ with affine transformation
$S_{0}^{\otimes w_{1}}\circ T_{0},S_{1}^{\otimes w_{1}},\ldots,S_{\ell}^{\otimes w_{1}},S_{0}^{\otimes w_{2}}\circ T_{1}\circ S_{\ell+1}^{\otimes w_{1}},S_{1}^{\otimes w_{2}},\ldots,S_{\ell}^{\otimes w_{2}},S_{0}^{\otimes w_{3}}\circ T_{2}\circ S_{\ell+1}^{\otimes w_{2}},\\ \ldots,\\ S_{0}^{\otimes w_{k}}\circ T_{k-1}\circ S_{\ell+1}^{\otimes w_{k-1}},S_{1}^{\otimes w_{k}},\ldots,S_{\ell}^{\otimes w_{k}},T_{k}\circ S_{\ell+1}^{\otimes w_{k}},$
we transform the NN with activation function $\tilde{\sigma}$ into a NN with activation function $\sigma$ with the desired characteristics. The claim about the size is immediate. Note that if $r=2$, then $\tilde{\sigma}$ is an affine map and the claim is still true. ∎